OpenStudy (anonymous):

MEDAL FOR THE PERSON THAT WRITES THE BEST ANSWER! I am having trouble with this problem can you help me solve it by showing me the steps on how to do it. It says: Kevin rolled a number cube (a die which is one of a pair of dice) 50 times and obtained the following results: Number which showed up : 1 , 2 , 3 , 4 , 5 , 6 Number of times it showed up : 10, 9 , 5 , 7 , 8 , 11 here are the questions: 1) what is the theoretical probability of rolling a "2"? 2) Based on these results, how many 5's should kevin expect to get in 300 rolls? 3) Based on theoretical probability, how many 5's should kevin expect to get in 300 rolls?

OpenStudy (ankit042):

1) Probability of rolling 2 = #times 2 came in the experiment/ total number of times experiment done 9/50 2) probability of 5 = 8/50 so for 300 rolls we will get 8/50*300= 48 3) In case of unbiased dice there is an equal chance for all the numbers so probability of 5 will be 1/6 and for 300 attempts we will get 1/6*300=50

OpenStudy (anonymous):

THANK YOU so much finally somebody answered and i really appreciate it .

OpenStudy (ankit042):

Did you understood the concept for the problem?

OpenStudy (anonymous):

yes i actually got two quesations correct on my own and i wanted to make sure i was right and thank you for the help! :)

OpenStudy (ankit042):

Great!